4. Do the following functions exhibit constant, increasing or decreasing returns to scale? a. l 31K2 b. Q L12K1 21 c.Q 4L1
Answer #1

0.0 4Lmk Now; Q - 4 (LM) (km) 2 4L2 mi2 kmi = mtl (4Lkk) To = mill (a so m3h.) ml : increasing return to Scale